# You Need to Pay Better Attention: Rethinking the Mathematics of Attention Mechanism

### v2

@@ 3.3 Super Attention: Introducing $W^{A}$
All three attention mechanisms, we discussed so far, have a learnable linear kernel between $Q$ and ${K}^{\intercal}$ but not between ${K}^{\intercal}$ and $V$. To better observe this, let us write the equation for one head in one of the attention mechanisms discussed so far, e.g., Efficient Attention by combining Equations (1416):
@@
As we see in Eq. 17, there are no learnable parameters between ${K}^{\intercal}$ and $V$, and the attention scores $S_{i}$ are directly applied to the values $V_{i}$. The intuition behind directly applying $S_{i}$ to $V_{i}$ is that the attention scores in $S_{i}$ determine “how much attention is paid” to each of the features of each token in $V_{i}$. Despite this intuition, we found out that by the absence of a learnable parameter here, the attention mechanism is performing suboptimally, in the sense that by introducing a new learnable parameter $W^{A}$ (following Principle 2), which sits between the attention scores $S_{i}$ and $V_{i}$, we can perform up to 17% better in vision and NLP tasks. Specifically, with the introduction of $W^{A}$, Eq. 17 evolves to
@@
The role of $W^{A}$ is to mix and align the values vertically (token-wise). Thus, to prevent “look ahead” in the attention mechanism for use in generative language modelling, we can constrain $W^{A}$ to be lower triangular, so that future tokens do not influence the current one in $W^{A}$. Note that we use the same $W^{A}$ for all heads. The reason here is that we want to improve the model performance while keeping the model size as small as possible. Thus, in a more general formulation, one can use different $W^{A}$ for each head to perhaps gain better performance, but at the cost of increasing the number of parameters, and thereby the model size.
Definition 4 (Super Attention).
Using the notation of Definition 3, Super Attention is the attention mechanism defined by the following set of equations:
@@ 4 Evaluation
We evaluate all the proposed mechanisms in vision (LABEL:subsec:Vision), NLP (Section 4.2). We also provided a detailed comparison of the computational costs and edge device performance in Sections B.1 and B.2, showcasing the efficiency and efficacy of the proposed attention mechanism on edge devices.
@@
We have chosen widely-used benchmarks that provide an easy and fair comparison between the four attention mechanisms used. In each experiment, where we have used a code from an online resource, we have minimised the changes to the original code to ensure the fairness of our comparison. For all experiments, (1) we use the same model architecture and iterate between standard, Optimised, Efficient, and Super Attention; (2) we use smallest model sizes that yield meaningful results to ensure maximum utilisation of the underlying attention mechanism; (3) we continue training until the validation loss flattens, which we have observed to be approximately the same among all attention mechanisms; and (4) we report the results by averaging over five runs (except for the ImageNet experiment, where due to the large dataset size, we report the results over one training run) to ensure the fairness of our comparison.
@@
All experiments are implemented in Keras with JAX backend using the examples available at keras.io/examples with minor dataset-specific adjustments, e.g., modifying the number of classes, layers, etc. All the reported results are obtained by training on an Nvidia RTX 4090 GPU (24GB VRAM) or an Nvidia A100 GPU (80GB VRAM); however, we have chosen model and batch sizes to ensure that they run on 24GB VRAM. In each table, we report the train and test accuracy and loss, the number of parameters in one attention layer (in the “# Param.” column), the average training time (in seconds) of models for one epoch on an RTX 4090 GPU (in the “Epoch Time” column), as well as other related task-specific metrics.
@@ 4.1 Vision Transformers
We use three widely-adopted vision dataset: MNIST(LeCun et al., 2010), CIFAR100 (Krizhevsky, 2009), and ImageNet1K Russakovsky et al. (2015). For Brevity, we refer to the ImageNet1K dataset throughout the paper as ImageNet. These datasets have a varying number of images with different sizes and different numbers of classes, thereby requiring different model sizes, which allows us to compare the proposed attention mechanisms for scalability.
Table 1: Averages of different metrics (over five runs on MNIST and CIFAR100, and one run on ImageNet). The numbers in parentheses indicate the ranking of each mechanism on each dataset for that metric. Ablation studies on the number of heads for all experiments is available in Section B.3. An additional ablation study for models of the same size but with different attention mechanisms is available in Section B.3. As expected, we observe that Efficient Attention models have the smallest attention layer size and the Super Attention models perform the best in terms of accuracy and loss.
Dataset Att. $h$ $d_{m}$ Param. Epoch Time Acc. (%) Loss Top 5 Test Acc. (%) Test Loss Test Top 5
@@
Stn. 8 256 263K (4) 21.19 (4) 50.38 (2) 1.85 (2) 80.85 (2) 34.64 (3) 2.89 (4) 63.69 (2)
Opt. 8 256 197K (2) 20.39 (3) 50.19 (3) 1.86 (3) 80.49 (3) 35.03 (2) 2.85 (2) 63.51 (3)
Eff. 8 256 131K (1) 19.22 (1) 48.96 (4) 1.90 (4) 79.98 (4) 34.54 (4) 2.85 (2) 63.43 (4)
CIFAR100 Sup. 8 256 197K (3) 20.28 (2) 58.68 (1) 1.50 (1) 86.51 (1) 35.97 (1) 2.84 (1) 64.29 (1)
Stn. 12 768 2.36M (4) 2572 (4) 51.18 (2) 2.09 (2) 76.05 (2) 32.74 (4) 3.36 (4) 56.48 (4)
Opt. 12 768 1.77M (3) 2426 (2) 50.98 (3) 2.09 (2) 75.76 (3) 36.31 (1) 3.09 (2) 60.65 (1)
Eff. 12 768 1.18M (1) 2374 (1) 49.42 (4) 2.14 (4) 74.94 (4) 36.17 (2) 3.07 (1) 60.39 (2)
ImageNet Sup. 12 768 1.22M (2) 2483 (3) 60.00 (1) 1.61 (1) 83.53 (1) 33.86 (3) 3.19 (3) 57.25 (3)
Analysis.
@@
The number of parameters in the models considered for the vision tasks range from 300K (MNIST) to 60M (ImageNet), their context length ranges from 64 (MNIST) to 256 (CIFAR100 and ImageNet), the dataset sizes range from 60K (MNIST) to 1.28M (ImageNet), and the number of classes ranges from 10 (MNIST) to 1K (ImageNet). We observe that in all experiments, Super Attention performs better than all other attention mechanisms. In particular, the performance gap between Super Attention and the rest increases with the model size and dataset complexity (up to $\nicefrac{{(60-51.18)}}{{51.18}}=17.23\%$ on ImageNet) despite having fewer parameters and being faster to train (up to $\nicefrac{{(8.31-7.58)}}{{7.58}}=5.86\%$ on MNIST). Efficient and Optimised Attention perform comparably to standard attention in terms of accuracy and loss but with $\nicefrac{{1}}{{2}}$ and $\nicefrac{{3}}{{4}}$ as many attention parameters, respectively. They are also significantly faster to train (up to $\nicefrac{{(8.31-7.05)}}{{8.31}}=15.16\%$ when comparing Efficient Attention to standard attention on MNIST).
4.2 Natural Language Processing
We evaluate the attention mechanisms introduced here in Transformer models of different sizes for two NLP tasks: sentiment classification and NMT, which was the main task in the seminal paper of Vaswani et al. (2017), “Attention is All You Need”. For sentiment classification, we use two widely-used benchmarks, IMDB Movie Reviews (Maas et al., 2011) and Amazon Reviews (Ni et al., 2019) datasets. For NMT, we use the combined Europarl (Koehn, 2005) and Anki (Anki.net, ) dataset for English-to-Spanish translation.
Table 2: Averages of different metrics over five runs in the natural language classification experiments on IMDB and Amazon Reviews datasets. The numbers in parentheses indicate the ranking of each mechanism for that metric on each dataset. Ablation studies on the number of heads for all experiments is available in LABEL:app:NLP. As expected, we observe that Efficient Attention models have the smallest attention layer size and the Super Attention models perform the best in terms of accuracy and loss.
Dataset Att. $h$ $d_{m}$ Param. Epoch Time Acc. (%) Loss Test Acc. (%) Test Loss
Stn. 4 32 4,224 (4) 0.315 (4) 95.70 (4) 0.086 (3) 77.62 (4) 0.474 (4)
Opt. 4 32 3,168 (2) 0.305 (3) 96.31 (3) 0.095 (4) 77.85 (2) 0.472 (2)
Eff. 4 32 2,112 (1) 0.280 (1) 96.41 (2) 0.064 (1) 77.77 (3) 0.468 (1)
IMDB Sup. 4 32 3,168 (2) 0.299 (2) 97.45 (1) 0.070 (2) 78.34 (1) 0.472 (2)
Stn. 4 64 16,640 (4) 20.38 (4) 62.54 (3) 0.868 (3) 52.74 (4) 1.097 (4)
Opt. 4 64 12,480 (2) 19.89 (3) 61.64 (4) 0.876 (4) 52.88 (3) 1.090 (3)
Eff. 4 64 8,320 (1) 17.20 (1) 63.55 (2) 0.845 (2) 53.19 (2) 1.080 (2)
Amazon Sup. 4 64 12,480 (2) 19.77 (2) 66.52 (1) 0.774 (1) 54.25 (1) 1.058 (1)
Table 3: Averages of different metrics over five runs for English-to-Spanish NMT on combined Europarl and Anki translation datasets. The numbers in parentheses indicate the ranking of each mechanism for that metric. Ablation studies on the number of heads for all experiments is available in Section B.4. As we observe, Optimised and Efficient Attentions perform similarly to standard attention on most metrics with $\nicefrac{{1}}{{2}}$ and $\nicefrac{{3}}{{4}}$ as many attention parameters, respectively. We did not include results for Super Attention, as it could not be directly compared against standard, Optimised, and Efficient Attention (see the analysis for details) in this task.
Att. $h$ $d_{m}$ $d_{k}$ Param. Epoch Time BLEU Acc. Loss Test BLEU Test Acc. Test Loss
Analysis.
The number of parameters in the models considered for the NLP tasks range from 650K (IMDB) to 2.6M (Amazon), their context length ranges from 32 (IMDB) to 64 (Amazon), the dataset sizes range from 50K (IMDB) to 3.65M (Amazon), and the number of classes ranges from 2 (IMDB) to 5 (Amazon). We observe a similar pattern to ViT for text classification in the sense that Super Attention significantly outperforms the other mechanisms in terms of train accuracy ($\nicefrac{{(66.52-62.64)}}{{62.64}}=6.19\%$ compared to standard attention on Amazon Reviews). We also observe that standard attention is the least efficient ($\nicefrac{{(600-523)}}{{523}}=14.72\%$ slower than Efficient Attention in NMT) with the highest number of parameters (twice as many parameters per layer compared to Efficient Attention). In the NMT experiment, we only include standard, Optimised and Efficient Attention because this task requires variable context length for the models. Variable context length for the Super Attention-based model could be achieved using techniques such as sliding-window attention (Beltagy et al., 2020), which would result in sparse attention and would not be directly comparable to the results in Table 3.
@@ 5 Related Work
After the adoption of Transformers, different research directions have emerged to address different shortcomings of the attention mechanism and Transformer models. Sparse attention, such as Longformer (Beltagy et al., 2020; Zhang et al., 2021a), reduces the computational complexity by focusing on key input parts (Child et al., 2019). Despite their efficiency in handling long sequences, sparse attention mechanisms struggle with tasks requiring comprehensive sequence analysis.
Flash Attention (Dao et al., 2022) and Flash Attention 2 (Dao, 2024) optimize multi-head attention for modern GPUs without changing its structure, enabling faster processing and reduced memory demands. It’s worth mentioning our proposed algorithms also benefit from these optimisations.
Quantization has revolutionised the adoption of FMs, particularly those based on Transformers. Recent advances include mixed-precision post-training quantization for vision transformers (Liu et al., 2021), quantization-aware training (Jacob et al., 2018; Nagel et al., 2022), mixed-precision training (Micikevicius et al., 2018), dynamic quantization (Zhang et al., 2021b), and layer-wise quantization (Chen et al., 2019). Moreover, Ding et al. (2022) unveiled a cutting-edge framework enhancing quantised model accuracy without significant performance degradation. However, quantization faces challenges such as potential performance drops and increased vulnerability to adversarial attacks (Hong et al., 2021; Gupta and Ajanthan, 2022).
Sparsification techniques reduce the number of non-zero elements in a network’s weights. Ashkboos et al. (2024) introduced a post-training sparsification technique for large language models that compresses weight matrices with 1-10% performance degradation. Increasing sparsity could lead to reduced robustness (Timpl et al., 2022).
6 Discussion
We hypothesised and validated, through mathematical reasoning and rigorous experiments, that the three proposed attention mechanisms, Optimised, Efficient, and Super Attention (see LABEL:fig:_AttentionFlowcharts for visualisations of the mechanisms), perform better or on par compared to standard attention while being respectively 25%, 50%, and 25% smaller in size and more computationally efficient.
Computer Vision.
We considered image classification on MNIST, CIFAR100, and ImageNet1K, comparing standard, Optimised, Efficient, and Super Attention. In terms of performance, i.e., accuracy and loss, Optimised and Efficient Attention performed similarly to standard attention, while having fewer parameters, and being faster to train and infer. Super Attention outperforms standard attention in terms of accuracy by 3.5%, 12.5%, and 17.25% on MNIST, CIFAR100, and ImageNet datasets, respectively, while being smaller and faster to train and infer.
Natural Language Processing.
We also considered sentiment classification on IMDB Movie Reviews and Amazon Reviews and NMT on combined Europarl Parallel Corpus and Anki datasets for English-to-Spanish translation. Optimised and Efficient Attention performed similarly to standard attention on all tasks while having fewer parameters and being faster, and Super Attention outperforms standard attention by 1.8% and 6.4% on IMDB and Amazon Reviews respectively.
Limitations.

### v3

@@ 3.3 Super Attention: Introducing $W^{A}$
Looking at the Eqs. (1-6), we observe that in SDPA, there are learnable parameters between $Q$ and $K$; however, there is no such parameter between $K$ and $V$ (even though a $\operatorname{softmax}$ is applied to the term containing $K$). Following Principle 2, we introduce a new learnable parameter $W^{A}$ that linearly transforms the values from the left. To better observe this, let us write the equation for one head in one of the attention variants, e.g., Efficient Attention by combining Eqs. (1416):
@@
As we see in Eq. 17, there are no learnable parameters between ${K}^{\intercal}$ and $V$, and the attention scores $S_{i}$ are directly applied to the values $V_{i}$. The intuition behind directly applying $S_{i}$ to $V_{i}$ is that the attention scores in $S_{i}$ determine “how much attention is paid” to each of the features of each token in $V_{i}$. Despite this intuition, we found that in practice the model can benefit from an additional kernel which comes in between the scores $S_{i}$ and values $V_{i}$. Specifically, with the introduction of $W^{A}$, Eq. 17 changes to
@@
The role of $W^{A}$ is to mix and align the values vertically (token-wise). Thus, to prevent “look ahead” in the attention mechanism for use in causal language modelling, we can constrain $W^{A}$ to be lower triangular, so that future tokens do not influence the current one in $W^{A}$. Note that we use the same $W^{A}$ for all heads. The reason here is that we want to improve the model performance while keeping the model size as small as possible. Thus, in a more general formulation, one can use different $W^{A}$ for each head to perhaps gain even better performance, but at the cost of increasing the number of parameters, and thereby the model size.
Definition 3.5 (Super Attention).
Using the notation of Def. 3.3, we define Super Attention with the following equations:
@@ 4 Evaluation
We evaluate the proposed mechanism on a range of NLP tasks (§ 4.1 and B.5); we then show that the approach generalises to other modalities by evaluating them on a number of vision benchmarks (LABEL:subsec:Vision). We also provided a detailed comparison of the computational costs and edge device performance in § 4.3, B.1, and B.2.
@@
We evaluate on a range of benchmarks. In each benchmark, we follow the common practices for evaluating the performances. For all benchmarks, (1) we use the same model architecture and iterate between standard, Optimized, Efficient, and Super Attention; (2) we continue training until validation loss flattens or a given computational budget is reached; and (3) for benchmarks on smaller datasets, we report the results by averaging over five runs to ensure fairness.
@@
All experiments in § 4.2 and 4.1 are implemented in Keras with JAX backend using keras.io/examples with minor dataset-specific adjustments, e.g., modifying the number of classes, layers, etc. The generative language modelling experiment in § 4.1 is an adaptation of Andrej Karpathy’s NanoGPT (Karpathy, 2022). All the reported results are obtained by training on an Nvidia RTX 4090 GPU (24GB VRAM) or an Nvidia A100 GPU (80GB VRAM); however, we have chosen model and batch sizes to ensure that they run on 24GB VRAM. In each table, we report the train and test loss and accuracy (where relevant), the number of parameters in one attention layer (in the “# Param.” column), the average training time (in seconds) of models for one epoch on an RTX 4090 GPU (in the “Epoch Time” column), as well as other related task-specific metrics.
4.1 NLP Benchmarks
In this section, we evaluate the attention variants in Transformer models of different scales for three NLP tasks: sentiment classification, Machine Translation (MT) and generative language modelling (LM) and NLI tasks.
Sentiment Classification.
For sentiment classification (Tbl. 1), we use two widely-used benchmarks, IMDB Movie Reviews (Maas et al., 2011) and Amazon Reviews (Ni et al., 2019) datasets. The dataset sizes for these two experiments in this part are 50k and 3.65M, and the model sizes are  650K and  26M parameters, respectively.
Machine Translation (MT)
For MT (Tbl. 2), we use the combined Europarl (Koehn, 2005) and Anki (Anki.net, ) dataset for English-to-Spanish translation. The dataset includes  2 million pairs and the model sizes range from  93-104 million parameters for different architectures.
Generative LM and NLI.
For generative language modelling (Tbl. 3), we use the OpenWebText dataset (Gokaslan and Cohen, 2019) for training and the HellaSwag dataset (Zellers et al., 2019) for comparing the common-sense reasoning performance of the trained models. This dataset includes more than 9 Billion tokens and the model sizes range between 110-124 million parameters for different architectures. The context window of the language models is set to 1024 tokens.
Table 1: Sentiment classification results, averaging over five runs on IMDB and Amazon Reviews datasets. Numbers in parentheses indicate the ranking of each attention variant for a given metric and dataset. Ablation studies on the number of heads for all experiments is available in LABEL:app:NLP. Efficient Attention models have the smallest attention layer size and the Super Attention models perform the best in terms of accuracy and loss.
Dataset Att. $h$ $d_{m}$ Param. Epoch Time Acc. (%) Loss Val Acc. (%) Val Loss
Stn. 4 32 4,224 (4) 0.315 (4) 95.70 (4) 0.086 (3) 77.62 (4) 0.474 (4)
Opt. 4 32 3,168 (2) 0.305 (3) 96.31 (3) 0.095 (4) 77.85 (2) 0.472 (2)
Eff. 4 32 2,112 (1) 0.280 (1) 96.41 (2) 0.064 (1) 77.77 (3) 0.468 (1)
IMDB Sup. 4 32 3,168 (2) 0.299 (2) 97.45 (1) 0.070 (2) 78.34 (1) 0.472 (2)
Stn. 4 128 66,048 (4) 66.97 (4) 88.49 (3) 0.25 (3) 65.55 (4) 0.77 (3)
Opt. 4 128 49,536 (3) 61.75 (3) 89.56 (1) 0.23 (1) 65.67 (2) 0.75 (2)
Eff. 4 128 33,024 (1) 56.44 (1) 86.63 (4) 0.29 (4) 65.58 (3) 0.77 (3)
Amazon Sup. 4 128 42,336 (2) 59.86 (2) 88.56 (2) 0.24 (2) 68.10 (1) 0.71 (1)
Table 2: Machine translation results, averaging over five runs for English-to-Spanish MT on combined Europarl and Anki translation datasets. Numbers in parentheses indicate the ranking of each attention variant for that metric. Ablation on the number of heads is available in § B.4. Optimized and Efficient Attentions perform similarly to standard attention on most metrics with $\nicefrac{{1}}{{2}}$ and $\nicefrac{{3}}{{4}}$ as many attention parameters, respectively. As the Super Attention layer has a fixed context length and the decoder requires a varying context length, using Super Attention would require using a sliding window, which would not be comparable to the full attention used for the other variants.
Att. $h$ $d_{m}$ $d_{k}$ Param. Epoch Time BLEU Acc. Loss Val BLEU Val Acc. Val Loss
NLP Results Analysis.
Super Attention outperforms attention variants in terms of validation accuracy (up to $\nicefrac{{(68.10-65.55)}}{{65.55}}=3.89\%$ compared to standard attention on Amazon Reviews) in the sentiment classification task. Similarly, we see for MT as well as generative LM and NLI tasks that the Optimized and Efficient architectures perform closely or on par with the Standard mechanisms. We also observe that standard attention is slower than all other variants (up to $\nicefrac{{(600-523)}}{{523}}=14.72\%$ slower than Efficient Attention in MT) with the highest number of parameters (twice as many parameters per layer compared to Efficient Attention). The generative LM experiment reveals subtle differences in performance among the models in training performance; However, our NLI experiment shows that when evaluated on the HellaSwag benchmark, all three models exhibit comparable performance, achieving accuracy rates between 30% and 31%.
Table 3: Averages of different metrics in generative LM using NanoGPT, a widely-referenced re-implementation of GPT-2 124M by Andrej Karpathy, based on different attention variants. The models are trained on the OpenWebText dataset ($\sim$9B training tokens) for one epoch with a batch size of 500 and a micro-batch size of 5 using a single A100 80GB node. The context window is 1024. In addition to the loss and perplexity, we provide the size of each model and the result of NLI on the HellaSwag benchmark. Similarly to the MT task, a fair comparison of Super Att. against other variants is not feasible as NanoGPT uses full attention but Super Att. requires using a sliding window.
Att. $h$ $d_{m}$ $d_{k}$ Layer Size Model Size Train Loss Train PPL Val Loss Val PPL HellaSwag
Stn. 12 768 64 2.36M 124M 2.92 18.5 3.13 22.9 0.31
Opt. 12 768 64 1.77M 117M 2.96 19.3 3.14 23.1 0.31
Eff. 12 768 64 1.18M 110M 3.02 20.5 3.18 24.0 0.30
@@ 4.2 Vision Transformers
We experiment with three widely adopted vision datasets of varying size and complexity: MNIST (LeCun et al., 2010), CIFAR100 (Krizhevsky, 2009), and ImageNet1K Russakovsky et al. (2015). For Brevity, we refer to the ImageNet1K dataset throughout the paper as ImageNet. Note that for the reported ImageNet results in Tbl. 4, we first pre-trained the model on the ImageNet21K dataset. We report the training details in § B.3.
Table 4: Vision results, averaging over five runs on MNIST and CIFAR100, and one run on ImageNet. Numbers in parentheses indicate the ranking of each mechanism for a given metric and dataset. An ablation study on the number of heads is available in § B.3. An additional ablation study for models of the same size on ImageNet but with different attention mechanisms is provided in § B.3. As expected, Efficient Attention models have the smallest attention layer size, and the Super Attention models achieve the highest accuracy and lowest loss.
Dataset Att. $h$ $d_{m}$ Param. Epoch Time Acc. (%) Loss Top 5 Val Acc. (%) Val Loss Val Top 5
@@
Stn. 8 256 263K (4) 21.19 (4) 72.28 (2) 1.41 (2) 91.02 (2) 48.14 (3) 1.82 (3) 90.22 (4)
Opt. 8 256 197K (2) 20.39 (3) 72.26 (3) 1.47 (3) 93.01 (3) 48.63 (2) 1.71 (2) 90.99 (2)
Eff. 8 256 131K (1) 19.22 (1) 71.96 (4) 1.49 (4) 92.23 (4) 47.95 (4) 1.83 (4) 90.48 (3)
CIFAR100 Sup. 8 256 197K (3) 20.28 (2) 79.62(1) 1.28 (1) 94.34 (1) 49.28 (1) 1.55 (1) 91.69 (1)
Stn. 12 768 2.36M (4) 2572 (4) 92.07 (2) 1.02 (2) 98.41 (2) 74.35 (3) 1.47 (3) 94.10 (4)
Opt. 12 768 1.77M (3) 2426 (2) 91.78 (3) 1.03 (3) 98.36 (3) 77.12 (2) 1.47 (3) 94.21 (3)
Eff. 12 768 1.18M (1) 2374 (1) 90.36 (4) 1.05 (4) 98.37 (4) 75.67 (4) 1.44 (2) 95.46 (2)
ImageNet Sup. 12 768 1.22M (2) 2483 (3) 94.09 (1) 0.94 (1) 99.32 (1) 79.29 (1) 1.39 (1) 96.37 (1)
ViT Results Analysis.
@@
The number of parameters in the models considered for the vision tasks range from 300K (MNIST) to 60M (ImageNet), their context length ranges from 64 (MNIST) to 256 (CIFAR100 and ImageNet), the dataset sizes range from 60K (MNIST) to 1.28M (ImageNet), and the number of classes ranges from 10 (MNIST) to 1K (ImageNet). Similar to text Transformers, ViTs using Super Attention architecture perform better than all other variants despite having fewer parameters than standard attention. Also, Optimized and Efficient Attentions perform comparably to standard attention with fewer parameters.
4.3 Speed and FLOPs Analysis
§ B.1 and B.2 are dedicated to studying the computational complexity and inference speed of the considered attention variants. Eq. 24 formulates the computational complexity for each algorithm.
Figs. 2 and 7 visualize a comparison between the required FLOPs for each algorithm based on “sequence length” and “projection dimension”. It indicates Efficient Attention requires the least number of FLOPs under all scenarios. From an empirical perspective, Tbl. 5 and Fig. 3 exhibit the faster inference speed (lower latency) of Efficient Attention compared to other variants in all datasets, followed by Optimized and Super variants.
(a) Standard Att.
(b) Efficient Att.
Figure 2: 3D plots visualizing the number of FLOPs for a forward + backward pass given different sequence lengths and projection dimensions in single-head setting for Efficient and Standard attention. Efficient Att. needs substantially fewer FLOPs for completing a forward + backward pass. Fig. 7 compares all architectures.
Figure 3: Summary of relative inference latency of models using different attention variants relative to standard attention on different datasets on Edge Device (Apple Laptop M2). Efficient Att. is the fastest (Optimized and Super Att. are also faster than standard attention). More details and numerical results are available in Tbl. 5.
Figure 4: Performance of different architectures on the Amazon Reviews as the size of models grows from 5 Million parameters to 25 Million parameters. In terms of test accuracy and loss, Super Attention shows increasingly better performance compared to all other architectures which are performing on par with each other. In terms of inference speed, all variants (especially Efficient) perform better than the Standard attention.
4.4 Scaling Analysis
We analyzed scaling behaviour across three dimensions: attention heads, dataset size, and model size. Head scaling experiments across tasks (Tbls. 5, 6, 8, 9 and 10) showed consistent performance improvements with increased heads for all architectures. Dataset scaling ranged from IMDB (50K examples) to OpenWebText (9B tokens) for language tasks, and MNIST (60K examples) to ImageNet (1.28M examples) for vision tasks, with our variants maintaining their relative performance advantages across scales. Model scaling experiments on Amazon Reviews (Figs. 4 and 8) demonstrate that as models grow from 5M to 25M parameters, Super Attention consistently outperforms standard attention, while Optimized and Efficient variants match standard’s performance with significantly fewer parameters. Notably, standard attention’s computational inefficiency becomes more pronounced at larger scales in both training and inference.
@@ 5 Related Work
Since their adoption, many research directions have emerged to address various shortcomings of attention mechanisms and Transformer models. Sparse attention, such as Longformer (Beltagy et al., 2020; Zhang et al., 2021a), reduces the computational complexity by focusing on key input parts (Child et al., 2019). Despite handling long sequences efficiently, sparse mechanisms struggle with tasks requiring a comprehensive sequence analysis.
Another line of research focuses on approximating the attention matrix to attain linear complexity. Performer (Choromanski et al., 2021) uses random feature maps and FAVOR+ mechanism; Linformer (Wang et al., 2020) projects keys and values to lower dimensions by exploiting low-rank properties. While these approaches achieve efficiency through approximation, they often compromise model quality. In contrast, our proposed variants achieve efficiency through structural modifications while maintaining or improving model quality.
Recent work has explored architectures that combine transformers’ parallel training capabilities with RNNs’ inference efficiency, including RWKV (Peng et al., 2023) with linear recurrence and State-Space models like Mamba (Gu and Dao, 2024) and S4 (Gu et al., 2021). While these approaches show promise, they require fundamental architectural changes. Our work instead focuses on optimizing the attention mechanism itself, preserving the proven benefits and versatility of transformer architectures while reducing computational costs.
Several approaches focus on reducing model redundancy. Voita et al. (2019) demonstrate that multi-head SDPA is over-parameterized, leading to collaborative frameworks that reduce projection sizes (Cordonnier et al., 2020). Similarly, sparsification techniques reduce non-zero elements in weights, with recent work achieving 1-10% compression with minimal performance impact (Ashkboos et al., 2024), though potentially affecting robustness (Timpl et al., 2022). While these approaches focus on post-hoc optimization or pruning, our work fundamentally reimagines the attention mechanism’s structure to achieve efficiency by design. We discuss further related attempts (including LoRA, Quantization and Flash Attention) for facilitating the deployability of transformers in § C.
6 Discussion and Conclusions
We proposed and evaluated three variants of SDPA that alter the standard arrangement of linear transformations to achieve better performance per computation cost and number of parameters (see Fig. 1 for visualizations). Optimized and Efficient Attention replace one (values) and two (values and keys) linear transformations with slicing, resulting in 25% and 50% size reductions and fewer matrix multiplications, respectively. The third variant, Super Attention, introduces a new linear transformation operating on the values from the left. While Super Attention can be applied to standard, Optimized, or Efficient Attention, we combined it with Efficient Attention, resulting in approximately 25% fewer parameters compared to standard attention.
Our evaluation spanned a wide range of tasks, including sentiment classification on IMDB and Amazon Reviews, Machine Translation on combined Europarl and Anki datasets, generative LM on OpenWebText dataset and NLI on HellaSwag. We used benchmarks varying in size from 50,000 examples to 9 billion tokens. To verify if these architectural benefits generalize across modalities, we also evaluated all variants for image classification on MNIST, CIFAR100, and ImageNet1K.
The experimental results demonstrate that Optimized and Efficient Attention performed comparably to standard attention across different benchmarks, despite having 25-50% fewer parameters and being faster. Super Attention consistently outperformed standard variant in all applicable benchmarks, achieving improvements of up to 10% on CIFAR100 and 4% on Amazon, while maintaining fewer parameters and faster training and inference.
Our generative LM experiment using a 1.1B Llama-based model in § B.5 provides insight into these variants’ performance at larger scales. Yet realizing their true potential requires evaluation at even larger scales, which are beyond our computational resources. The promising results suggest these attention variants could open new pathways for training and deploying capable models on devices with limited computational resources, like smartphones and small personal devices.
Limitations